OpenStudy (anonymous):

Evaluate the indefinite integral. 1/cos^5(x)dx

OpenStudy (anonymous):

sounds like a calc II problem to me

OpenStudy (anonymous):

yeah its a calc 2 problem....its giving me trouble.

myininaya (myininaya):

\[\int\limits_{}^{}\sec^5(x) dx=\int\limits_{}^{}\sec^4(x) \sec(x) dx=\int\limits_{}^{}(1+\tan^2(x))^2\sec(x) dx\] \[=\int\limits_{}^{}(1+2\tan^2(x)+\tan^4(x))\sec(x) dx\] \[=\int\limits_{}^{}\sec(x) dx+2 \int\limits_{}^{}\tan^2(x) \sec(x) dx+\int\limits_{}^{}\tan^4(x) \sec(x) dx\] ---------------------------------------------------------------------- first lets do \[\int\limits_{}^{}\sec(x) dx=\int\limits_{}^{}\sec(x) \cdot \frac{\sec(x)+\tan(x)}{\sec(x)+\tan(x)} dx\] let \[u=\sec(x)+\tan(x) => du=(\tan(x)\sec(x)+\sec^2(x) )dx=\sec(x)(\tan(x)+\sec(x))dx\] so we have \[\int\limits_{}^{}\sec(x) dx=\int\limits_{}^{}\sec(x) \cdot \frac{\sec(x)+\tan(x)}{\sec(x)+\tan(x)} dx\] \[=\int\limits_{}^{}\frac{du}{u}=\ln|u|+c_1=\ln|\sec(x)+\tan(x)|+c_1\] ----------------------------------------------------------------------.....

myininaya (myininaya):

now we look at \[2\int\limits_{}^{}\tan^2(x) \sec(x) dx\]

OpenStudy (anonymous):

reminds me of why i hate this subject so very much

myininaya (myininaya):

let's look at \[\int\limits_{}^{}\tan^2(x)\sec(x) dx\]

myininaya (myininaya):

and then when we find antiderivative we will multiply by 2

myininaya (myininaya):

\[\int\limits_{}^{}\tan^2(x)\sec(x) dx=\int\limits_{}^{}(\sec^2(x)-1)\sec(x) dx=\int\limits_{}^{}\sec^3(x) dx-\int\limits_{}^{}\sec(x) dx\]

OpenStudy (anonymous):

so you have to integral by parts the sec^2 and sec^3

myininaya (myininaya):

but we know \[\int\limits_{}^{}\sec^3(x) dx-\int\limits_{}^{}\sec(x) dx=\int\limits_{}^{}\sec^3(x) dx-\ln|\sec(x)+\tan(x)| +c_2\] lol i love my constants anyways now we to do \[\int\limits_{}^{}\sec^3(x) dx=\int\limits_{}^{}\sec^2(x)\sec(x) dx\] we need integration by parts here

myininaya (myininaya):

integral by parts?

myininaya (myininaya):

i just use the distributive property

OpenStudy (anonymous):

lol i mean integrate by part

myininaya (myininaya):

\[\int\limits_{}^{}\sec^2(x) \sec(x) dx=\tan(x) \sec(x)-\int\limits_{}^{}\tan(x) \sec(x) \tan(x) dx\] \[=\tan(x) \sec(x)-\int\limits_{}^{}\tan^2(x) \sec(x) dx\] \[=\tan(x)\sec(x)-\int\limits_{}^{}(\sec^2(x)-1)\sec(x) dx\] \[=\tan(x)\sec(x)-\int\limits_{}^{}\sec^3(x) dx+\int\limits_{}^{}\sec(x) dx\] \[=\tan(x)\sec(x)-\int\limits_{}^{}\sec^3(x) dx+\ln|\sec(x)+\tan(x)|\] add \[\int\limits_{}^{}\sec^3(x) dx\] on both sides then divide by 2 to get \[\int\limits_{}^{}\sec^3(x) dx=\frac{1}{2}\tan(x)\sec(x)+\frac{1}{2}\ln|\sec(x)+\tan(x)|\]

myininaya (myininaya):

\[\int\limits\limits_{}^{}\tan^2(x)\sec(x) dx=\int\limits\limits_{}^{}(\sec^2(x)-1)\sec(x) dx=\int\limits\limits_{}^{}\sec^3(x) dx-\int\limits\limits_{}^{}\sec(x) dx \] \[=\frac{1}{2}\tan(x)\sec(x)+\frac{1}{2}\ln|\sec(x)+\tan(x)|-\ln|\sec(x)+\tan(x)|+c_3\] so anyways don't forget we actually had \[2\int\limits_{}^{}\tan^2(x)\sec(x) dx=2 \cdot (\frac{1}{2} \tan(x)\sec(x)+\frac{1}{2}\ln|\sec(x)+\tan(x)|-\ln|\sec(x)+\tan(x)|)+c_3\]

myininaya (myininaya):

\[2\int\limits_{}^{}\tan^2(x)\sec(x) dx=\tan(x)\sec(x)-\ln|\sec(x)+\tan(x)|+c_3\]

myininaya (myininaya):

ok then the last part is \[\int\limits_{}^{}\tan^4(x) \sec(x) dx\] i leave this to you as an exercise

OpenStudy (anonymous):

okay thanks you for showing me the steps on how i can reach this problem and ones similar to it